Find the equation of a line with-in y-intercept (0,4) and m=1/3
Answer by algebrahouse.com(1659)

slope = 1/3
y-intercept = 4

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

y = (1/3)x + 4 {substituted into slope-intercept form}

For more help from me, visit: www.algebrahouse.com